OpenStudy (anonymous):

Four out of 9 performers will be chosen to stand in a row on stage. How many ways can the 4 performers stand in a row? 362,880 ways 347,760 ways 15,120 ways 3024 ways

OpenStudy (paki):

5*6*7*8*9 = 15120 ways...

OpenStudy (irishboy123):

The answer they want is calculated as follows: \[nPr = \frac{ n! }{ (n-r)! }\] n = 9, r = 4 the real answer is 12! read the question again, it is very badly written.